# TF-Restormer: Complex Spectral Prediction for Speech Restoration

### v2

@@ 4 Training
The model is trained by two phases of pretraining and adversarial training. The model is trained with $f_{D}$ randomly selected from {16, 24, 44.1, 48}kHz at each step by downsampling target speech signals from VCTK dataset (Yamagishi et al., 2019). Based on speech sources, we simulated noisy reverberant signals by convolving the room impulse response (RIR) and noise samples from the DNS dataset (Reddy et al., 2020). We then applied various digital distortions including codecs and downsampled the signal to the sampling rates $f_{E}$ of 8k or 16kHz, which are common in practical restoration condition (see Appendix B.1 for details).
@@ 4.1 Pretraining
Following Babaev et al. (2024), we incorporate a self-supervised learning (SSL)-based perceptual loss to stabilize adversarial training and encourage human-aligned quality. Specifically, extracting features from a pretrained SSL model for both the enhanced and clean waveforms, we minimize the mean-squared-error between these representations:
$\mathcal{L}_{\text{p}}(\theta)=\mathbb{E}_{m,n}\!\left[\,|\phi(g_{\theta}(x))_{m,n}-\phi(s)_{m,n}|^{2}\right],$ (3)
given that ${y}=g_{\theta}(x)$ is output of restoration model $g_{\theta}(\cdot)$ with parameters of $\theta$. $\phi(\cdot)_{m,n}$ denotes the $m$-th element of $n$-th frame from its feature map. We utilize WavLM-conv (Chen et al., 2022) as in the previous study (Babaev et al., 2024).
@@
where $w_{tf}$ is scale factor that controls the relative scaling of gradient. The formulation $w\log(1+d/w)$ formulation ensures large gradients on smaller distance $d$ than $w$, preserving regions where phase and magnitude are already reconstructed well, while suppressing the influence of large deviations, avoiding the averaging common in $\ell_{1}$ or $\ell_{2}$ losses. For choosing the weight value $w_{tf}$, we observed that a distance $|Y_{c,tf}-S_{c,tf}|$ tends to be proportional to the scale of the corresponding source component $S_{m,tf}$. Therefore, $w_{tf}$ is empirically set to $E[S_{m,tf}]$ by averaging over the frames. We use $\alpha_{m}\hskip-1.42262pt=\hskip-1.42262pt0.6$, $\alpha_{r}\hskip-1.42262pt=\hskip-1.42262pt0.2$, and $\alpha_{i}\hskip-1.42262pt=\hskip-1.42262pt0.2$. Finally, when combined with the perceptual loss, our proposed objective becomes
$\mathcal{L}_{\rm pre}(\theta)=\lambda_{\rm p}\mathcal{L}_{\rm p}(\theta)+\lambda_{\rm s}\mathcal{L}_{\rm s}(\theta).$ (5)
where $\lambda_{*}$ denote loss weighting factors. Therefore, perceptual loss mainly focuses on largely deviated component while the scaled log-spectral more on well-predicted component, making them complementary to each other. For pretraining, we train the TF-Restormer with $\lambda_{\text{p}}\hskip-1.42262pt=\hskip-1.42262pt100$ and $\lambda_{\text{s}}\hskip-1.42262pt=\hskip-1.42262pt1$.
@@ 4.2 Adversarial Training
$\mathcal{L}_{\rm gen}(\theta)\hskip 7.11317pt$ $=$ $\lambda_{\rm g}\mathcal{L}_{\rm g}(\theta)+\lambda_{\rm fm}\mathcal{L}_{\rm fm}(\theta)+\lambda_{\rm p}\mathcal{L}_{\rm p}(\theta)+\lambda_{\rm s}\mathcal{L}_{\rm s}(\theta)+\lambda_{\rm hf}\mathcal{L}_{\rm hf}(\theta),$ (6)
$\mathcal{L}_{\rm disc}(\varphi_{i})$ $=$ $\mathcal{L}_{\rm d}(\varphi_{i}),\quad i=1,...,I.$ (7)
where $\mathcal{L}_{\text{hf}}=\mathcal{L}_{\text{pesq}}+10\cdot\mathcal{L}_{\text{utmos}}$ is additional human-feedback loss (Babaev et al., 2024) for aesthetic quality with differentiable PESQ loss [^1] https://github.com/audiolabs/torch-pesq and UTMOS loss (Saeki et al., 2022). We performed adversarial training using $\mathcal{L}_{\text{gen}}(\theta)$ with $\lambda_{\text{g}}\hskip-1.42262pt=\hskip-1.42262pt0.005$, $\lambda_{\text{fm}}\hskip-1.42262pt=\hskip-1.42262pt0.1,\lambda_{\text{p}}\hskip-1.42262pt=\hskip-1.42262pt100$, $\lambda_{\text{s}}\hskip-1.42262pt=\hskip-1.42262pt1$, and $\lambda_{\text{hf}}\hskip-1.42262pt=\hskip-1.42262pt0.0001$. Notably, we assign small weights to $\mathcal{L}_{\text{g}}$ and $\mathcal{L}_{\text{hf}}$ to avoid excessive generation artifacts. (See Appendix B.2 for details of training configuration.)
@@
In conventional adversarial training, a dedicated generator for each target sampling rate is trained with a corresponding discriminator as well (Défossez et al., 2023; Babaev et al., 2024; Ju et al., 2024). For stable training of a single generator across diverse rates, we propose a discriminator based on SFI-STFT, which preserves a consistent physical frame duration across sampling rates. Implemented with strided Conv2D layers, STFT discriminator (Défossez et al., 2023) produces two-dimensional maps that provide local real/fake supervision in the time-frequency plane. This design maintains sensitivity to spectral structure while remaining agnostic to absolute frequency resolution, thereby supporting adversarial training across different rates without redundant resampling or multiple discriminators. We employ 5 SFI discriminators with STFT window durations of {20, 40, 60, 80, 100} ms.
@@ 5 Evaluation
5.1 Dataset and Metrics
UNIVERSE data for speech restoration
As speech restoration model, we evaluate on 100 synthetic samples generated by UNIVERSE authors (Serrà et al., 2022) to ensure comparability to prior works in $f_{E}=f_{D}=16$kHz setting. The dataset introduces diverse simulated degradations such as bandpass filtering, reverberation, codec compression, and transmission artifacts.
VCTK-DEMAND for denoising
We additionally evaluated the well-known Valentini denoising dataset (Valentini-Botinhao et al., 2017) for direct comparison with conventional enhancement models as speech enhancement benchmarking. The evaluation set (824 utterances) consists of noisy mixtures from two speakers under four SNR conditions (17.5, 12.5, 7.5, and 2.5 dB).
VCTK for super-resolution
For super-resolution evaluation, we construct paired data by downsampling 48 kHz clean utterances from the VCTK-0.92 dataset (Yamagishi et al., 2019). Beyond the clean case, we also create noisy-distorted conditions by adding degradations such as noise, reverberation, band-pass filtering, and codec effects, enabling a comprehensive evaluation of speech restoration with super-resolution. Note that the training simulation follows a similar procedure, which may provide a slight advantage to our model. (See Appendix D for details of the simulation.)
For the evaluation, we adopt non-intrusive perceptual estimators for mean opinion score (MOS): DNSMOS (Reddy et al., 2022), UTMOS (Saeki et al., 2022), and WVMOS (Andreev et al., 2022) to assess the perceptual quality regardless of reference signals. We also employ perceptual evaluation of speech quality (PESQ) (Rix et al., 2001) to assess the perceptual signal fidelity of restored signal compared to the reference. As a complement of PESQ based on lower band, we consider signal-to-distortion ratio (SDR) (Roux et al., 2018), log-spectral distance (LSD), and mel-cepstral distortion (MCD) (Fukada et al., 1992) to assess the full-band signal (44.1/48kHz). In addition, to evaluate the reference-aware speech generation quality by capturing semantic congruence, we report SpeechBERTScore(sBERT) and SpeechTokenDistance(sTokDis) (Saeki et al., 2024). We also confirm effectiveness on real recordings based on DNSMOS, UTMOS, and WVMOS (Appendix F).
5.2 Comparison with existing models
For the UNIVERSE dataset, we consider VoiceFixer (Liu et al., 2022b) as a Mel vocoder-based baseline, StoRM (Lemercier et al., 2023), UNIVERSE (Serrà et al., 2022), and UNIVERSE++ (Scheibler et al., 2024) as diffusion-based baselines, TF-Locoformer as a recent TF dual-path Transformer model, and FINALLY (Babaev et al., 2024) as a latest strong Mel-vocoder method. As shown in Table 1, VoiceFixer improves MOS but sacrifices fidelity due to its Mel representation, while FINALLY achieves the highest perceptual quality yet lacks signal fidelity, a trend confirmed in Table 2. Diffusion-based methods yield more balanced results by directly operating in the waveform or complex STFT. TF-Locoformer preserves signal-level fidelity but suffers from residual perceptual artifacts and failure to recover lost details and naturalness (MOS, LSD). In contrast, TF-Restormer provides consistent improvements under fidelity, semantics, and perceptual quality, with its streaming variant maintaining competitive effectiveness under causal constraints. This indicates its robustness across diverse degradations in a universal restoration setting. Note that all the compared models are offline methods.
@@ 5.2 Comparison with existing models
Model Signal fidelity Semantic fidelity Non-intrusive quality
[…]
Input 1.55 5.58 1.89 10.21 0.84 0.69 1.76 2.19 2.23
Ground Truth 4.50 $\infty$ 0.00 0.00 1.00 1.00 4.28 4.26 3.33
VoiceFixer 1.77 -5.68 1.49 10.50 0.84 0.71 3.28 2.83 2.99
StoRM 1.76 9.01 1.67 6.87 0.84 0.70 3.14 2.70 2.94
@@
TF-Restormer 2.30 11.12 1.45 5.08 0.91 0.80 4.34 4.08 3.25
TF-Restormer-streaming 2.00 8.89 1.47 6.01 0.87 0.74 3.93 3.77 3.14
Next, we evaluate TF-Restormer on the VCTK+DEMAND focusing on denoising. In Table 2, we compare against DB-AIAT (Yu et al., 2021), MP-SENet (Lu et al., 2023), and TF-Locoformer as dedicated denoising models, and VoiceFixer, UNIVERSE, and FINALLY as universal restoration baselines. Since the input speech is already well preserved and only corrupted by additive noise, it favors models that minimize unnecessary generation and faithfully retain the input signal. Accordingly, dedicated denoising models outperform universal restoration models in terms of signal fidelity, as they are optimized to suppress noise without altering intact regions. In contrast, restoration models risk degrading reliability by over-modifying clean inputs, making them less trustworthy for such simple cases. While not surpassing dedicated denoising models in raw signal metrics, TF-Restormer achieves more consistent semantic and perceptual gains, showing strong generalization despite being designed for universal restoration.
Table 2: Results on VCTK-DEMAND for denoising task.
Model Signal fidelity Semantic fidelity Non-intrusive quality
@@
Input 1.98  8.56 1.27 5.40 0.91 0.82 3.01 2.90 2.45
Ground Truth 4.50 $\infty$ 0.00 0.00 1.00 1.00 4.52 4.07 3.16
DB-AIAT 3.27 21.30 0.90 1.77 0.95 0.87 4.39 3.83 3.13
MP-SENet 3.61 21.03 0.85 1.58 0.95 0.88 4.35 3.86 3.12
TF-Locoformer 3.30 23.82 0.92 3.58 0.95 0.87 4.66 3.93 3.20
VoiceFixer 2.40 -1.12 0.97 7.40 0.90 0.81 4.15 3.50 3.08
UNIVERSE 2.84 18.77 1.17 2.20 0.92 0.83 4.32 3.75 3.03
FINALLY 2.94   4.60 - - - - 4.87 4.32 3.22
TF-Restormer 3.41 19.45 0.75 1.54 0.95 0.88 4.75 4.14 3.14
TF-Restormer-streaming 2.89 16.43 0.85 2.16 0.93 0.84 4.56 4.05 3.09
Finally, we experiment on the super-resolution task in Table 3, using a single model that directly supports arbitrary output sampling rates. For clean cases, we compare against dedicated super-resolution models: NVSR (Liu et al., 2022a), Frepainter (Kim et al., 2024), and AP-BWE (Lu et al., 2025), as well as VoiceFixer as a universal restoration baseline. As in Table 2, since the low-band of the input speech remains intact, dedicated models that concentrate on reconstructing the upper bands are favored. Unlike conventional approaches that rely on fixed input-output rates and often require zero-padding or redundant resampling, TF-Restormer leverages extension queries to dynamically expand the spectrum. With this versatility, TF-Restormer shows stable performance comparable to the dedicated models, faithfully retaining clean low-frequency regions while effectively generating high-frequency components. In adittion, under noisy-distorted conditions, TF-Restormer simultaneously restores corrupted regions and reconstructs missing high bands, demonstrating robust generalization beyond pure super-resolution. Overall, these results suggest the advantage of our model as a universal restoration framework that achieves bandwidth extension without sacrificing signal fidelity or requiring explicit resampling.
@@
Method 8kHz $\rightarrow$ 16kHz 8kHz $\rightarrow$ 24kHz 8kHz $\rightarrow$ 44.1kHz 16kHz $\rightarrow$ 48kHz
@@
clean (super-resolution only)
Input 2.53 1.84 3.33 2.91 2.03 3.33 3.44 2.44 3.33 3.17 1.31 3.83
@@
TF-Restormer 0.89 1.29 4.10 0.95 1.48 4.09 1.01 1.74 4.10 0.97 1.28 4.11
Noisy-distorted (general speech restoration + super-resolution)
Input 3.36 11.38 1.73 3.49 11.60 1.73 3.64 11.47 1.73 3.48 11.37 1.73
@@
StoRM 1.76 4.57 2.57 - - - - - - - - -
UNIVERSE++ 1.79 5.28 2.49 - - - - - - - - -
TF-Restormer 1.16 2.78 4.22 1.21 2.97 4.22 1.18 3.08 4.21 1.18 2.86 4.22
TF-Restormer-streaming 1.30 3.93 3.96 1.31 4.01 3.99 1.30 4.05 4.02 1.26 3.86 4.05
Table 4: Ablation Study. VCTK-C and -ND denote clean and noisy-distorted input from VCTK data in Table 3.
@@
UNIVERSE (general restoration)
Mag.-only 2.07 6.03 3.82
Mag.+Cplx. 2.23 5.70 3.76
Scaled log 2.29 4.96 4.10
VCTK+DEMAND (denoising)
Mag.-only 2.93 3.13 3.95
Mag.+Cplx. 2.97 2.98 3.87
Scaled log 3.41 1.54 4.14
VCTK-C (super-resol., 8 $\to$ 16kHz)
Mag.-only 3.42 2.10 4.10
Mag.+Cplx. 3.48 1.86 4.07
Scaled log 3.70 1.29 4.10
(a) Effect of scaled log-spectral loss
[…]
VCTK-C (super-resol., 8 $\to$ 16kHz)
Dec-only. 2.12 2.93 3.82
w/o MHCA 1.04 1.81 4.07
w/ MHCA 0.89 1.29 4.10
VCTK-C (super-resol., 8 $\to$ 44.1kHz)
Dec-only. 3.25 10.48 3.36
w/o MHCA 1.35 1.70 4.07
w/ MHCA 1.01 1.74 4.10
VCTK-ND (super-resol., 8 $\to$ 16kHz)
Dec-only. 2.23 4.39 3.77
w/o MHCA 1.20 3.30 4.11
w/ MHCA 1.16 2.78 4.22
(b) Decoder design
[…]
UNIVERSE (general restoration)
w/o F-proj. 2.26 5.56 3.90
w/ F-proj. 2.29 4.96 4.10
VCTK+DEMAND (denoising)
w/o F-proj. 3.21 0.85 4.03
w/ F-proj. 3.41 0.75 4.14
VCTK-C (super-resol., 8 $\to$ 16kHz)
w/o F-proj. 3.54 1.68 3.97
w/ F-proj. 3.70 1.29 4.11
(c) Effects of frequency projection

### v3

@@ 4 Training
The model is trained by two phases of pretraining and adversarial training. The model is trained with $f_{D}$ randomly selected from {16, 24, 44.1, 48}kHz at each step by downsampling target speech signals from VCTK dataset (Yamagishi et al., 2019). Based on speech sources, we simulated noisy reverberant signals by convolving the room impulse response (RIR) and noise samples from the DNS dataset (Reddy et al., 2020). We then applied various digital distortions including codecs and downsampled the signal to the sampling rates $f_{E}$ of 8k or 16kHz, which are common in practical restoration condition (see Appendix A.1 for details). Because extension query could be undertrained if distribution of the input and output sample rates is unbalanced in training, we investigate these issue in Appendix E.
@@ 4.1 Pretraining
Following (Babaev et al., 2024), we incorporate a self-supervised learning (SSL)-based perceptual loss to stabilize adversarial training and encourage human-aligned quality. Specifically, extracting features from a pretrained SSL model for both the enhanced and clean waveforms, we minimize the mean-squared-error between these representations:
$\mathcal{L}_{\text{p}}(\theta)=\mathbb{E}_{m,n}\!\left[\,|\phi(g_{\theta}(x))_{m,n}-\phi(s)_{m,n}|^{2}\right],\vskip-1.42262pt$ (3)
given that ${y}=g_{\theta}(x)$ is output of restoration model $g_{\theta}(\cdot)$ with parameters of $\theta$. $\phi(\cdot)_{m,n}$ denotes the $m$-th element of $n$-th frame from its feature map. We utilize WavLM-conv (Chen et al., 2022b) as in the previous study (Babaev et al., 2024).
@@
where $w_{tf}$ is scale factor that controls the relative scaling of gradient. The formulation $w\log(1+d/w)$ ensures large gradients on smaller distance $d$ than $w$, preserving regions where phase and magnitude are already reconstructed well, while suppressing the influence of large deviations, avoiding the averaging common in $\ell_{1}$ or $\ell_{2}$ losses. For choosing the weight value $w_{tf}$, we empirically observed that a distance $|Y_{c,tf}-S_{c,tf}|$ tends to be proportional to the source $|S_{m,tf}|$. Therefore, $w_{tf}$ is set to $E[S_{m,tf}]$ by averaging over the frames. We use $\alpha_{m}\hskip-1.42262pt=\hskip-1.42262pt0.6$, $\alpha_{r}\hskip-1.42262pt=\hskip-1.42262pt0.2$, and $\alpha_{i}\hskip-1.42262pt=\hskip-1.42262pt0.2$. Finally, when combined with the perceptual loss, our proposed objective becomes $\mathcal{L}_{\rm pre}(\theta)=\lambda_{\rm p}\mathcal{L}_{\rm p}(\theta)+\lambda_{\rm s}\mathcal{L}_{\rm s}(\theta)$ where $\lambda_{*}$ denote loss weighting factors. Therefore, perceptual loss mainly focuses on largely deviated component while the scaled log-spectral more on well-predicted component, making them complementary to each other. For pretraining, we train the TF-Restormer with $\lambda_{\text{p}}\hskip-1.42262pt=\hskip-1.42262pt100$ and $\lambda_{\text{s}}\hskip-1.42262pt=\hskip-1.42262pt1$.
@@ 4.2 Adversarial training
$\hskip-17.07164pt\mathcal{L}_{\rm gen}(\theta)\hskip-2.84526pt$ $=$ $\hskip-5.69054pt\lambda_{\rm g}\mathcal{L}_{\rm g}(\theta)+\lambda_{\rm fm}\mathcal{L}_{\rm fm}(\theta)$ (5)
$\hskip 14.22636pt+\lambda_{\rm p}\mathcal{L}_{\rm p}(\theta)+\lambda_{\rm s}\mathcal{L}_{\rm s}(\theta)+\lambda_{\rm hf}\mathcal{L}_{\rm hf}(\theta),$
$\hskip-17.07164pt\mathcal{L}_{\rm disc}(\varphi_{i})\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\mathcal{L}_{\rm d}(\varphi_{i}),\quad i=1,...,I.\vskip-2.84526pt$ (6)
where $\mathcal{L}_{\text{hf}}=\mathcal{L}_{\text{pesq}}+10\cdot\mathcal{L}_{\text{utmos}}$ is additional human-feedback loss (Babaev et al., 2024) for aesthetic quality with differentiable PESQ loss and UTMOS loss (Saeki et al., 2022). We performed adversarial training using $\mathcal{L}_{\text{gen}}(\theta)$ with $\lambda_{\text{g}}\hskip-1.42262pt=\hskip-1.42262pt0.005$, $\lambda_{\text{fm}}\hskip-1.42262pt=\hskip-1.42262pt0.1,\lambda_{\text{p}}\hskip-1.42262pt=\hskip-1.42262pt100$, $\lambda_{\text{s}}\hskip-1.42262pt=\hskip-1.42262pt1$, and $\lambda_{\text{hf}}\hskip-1.42262pt=\hskip-1.42262pt0.0001$. Notably, we assign small weights to $\mathcal{L}_{\text{g}}$ and $\mathcal{L}_{\text{hf}}$ to avoid excessive generation artifacts.
@@
In conventional adversarial training, a dedicated generator for each target sampling rate is trained with a corresponding discriminator as well (Défossez et al., 2023; Babaev et al., 2024; Ju et al., 2024), which introduces implementation overhead depending on the output sample rates. For training of a single generator across diverse rates, we propose a discriminator based on SFI-STFT, which preserves a consistent physical frame duration across sampling rates. Implemented with strided Conv2D layers, STFT discriminator (Défossez et al., 2023) produces two-dimensional maps that provide local real/fake supervision in the time-frequency plane. This design maintains sensitivity to spectral structure while remaining agnostic to absolute frequency resolution, thereby supporting adversarial training across different rates without redundant resampling or multiple discriminators. We employ 5 SFI discriminators with STFT window durations of {20, 40, 60, 80, 100} ms.
@@ 5 Evaluation
5.1 Test dataset and metrics
We evaluate a unified TF-Restormer for various datasets to validate the robustness across heterogeneous distortion conditions and arbitrary input–output sample-rates. For denoising(DN) and speech super-resolution (SSR), we also reports dedicated version of TF-Restormer for fair comparison and training configuration for dedicated models are summarized in Appendix A.3.
UNIVERSE data for general speech restoration (GSR)
As GSR model, we evaluate on 100 synthetic samples generated by UNIVERSE authors (Serrà et al., 2022) to ensure comparability to prior works in $f_{E}=f_{D}=16$kHz setting. The dataset introduces diverse simulated degradations such as bandpass filtering, reverberation, codec compression, and transmission artifacts.
VCTK-DEMAND for DN
We additionally evaluated the well-known Valentini denoising dataset (Valentini-Botinhao and others, 2017) for direct comparison with conventional enhancement models as speech enhancement benchmarking. The evaluation set (824 utterances) consists of noisy mixtures from two speakers under four SNR conditions (17.5, 12.5, 7.5, and 2.5 dB).
VCTK for SSR
For SSR evaluation, we construct paired data by downsampling 48 kHz clean utterances from the VCTK-0.92 dataset (Yamagishi et al., 2019). Beyond the clean case, we also create noisy-distorted conditions by adding degradations such as noise, reverberation, band-pass filtering, and codec effects, enabling a comprehensive evaluation of GSR with SSR. Note that the training simulation follows a similar procedure, which may provide a slight advantage to our model.
For the evaluation, we adopt non-intrusive perceptual estimators for mean opinion score (MOS): DNSMOS (Reddy et al., 2022), UTMOS (Saeki et al., 2022), and NISQA (Mittag et al., 2021) to assess the perceptual quality. Also, to assess the perceptual signal fidelity of restored signal compared to the reference, we employ perceptual evaluation of speech quality (PESQ) (Rix et al., 2001), signal-to-distortion ratio (SDR) (Le Roux et al., 2019), log-spectral distance (LSD), mel-cepstral distortion (MCD) (Fukada et al., 1992). In addition, to evaluate the reference-aware speech generation quality by capturing semantic congruence, we report SpeechBERTScore(sBERT) (Saeki et al., 2024).
@@ 5.1 Test dataset and metrics
Model Signal fidelity Perceptual quality
PESQ SDR LSD MCD sBERT UTMOS DNSMOS
Input 1.55 5.58 1.89 10.21 0.84 2.19 2.23
Ground Truth 4.50 $\infty$ 0.00 0.00 1.00 4.26 3.33
VoiceFixer 1.77 -5.68 1.49 10.50 0.84 2.83 2.99
StoRM 1.76 9.01 1.67 6.87 0.84 2.70 2.94
@@
TF-Restormer (off) 2.30 11.12 1.45 5.08 0.91 4.08 3.25
TF-Restormer (on) 2.00 8.89 1.47 6.01 0.87 3.77 3.14
5.2 Validation of stability across diverse scenarios
In this section, we validate the operational stability of TF-Restormer across a wide spectrum of restoration scenarios. For the UNIVERSE dataset, we consider VoiceFixer (Liu et al., 2022b) as a Mel vocoder-based baseline, StoRM (Lemercier et al., 2023), UNIVERSE (Serrà et al., 2022), and UNIVERSE++ (Scheibler et al., 2024) as diffusion-based baselines, TF-Locoformer as a recent TF dual-path Transformer model, and FINALLY (Babaev et al., 2024) as a latest strong Mel-vocoder method. As shown in Table 1, VoiceFixer improves MOS but sacrifices fidelity due to its Mel representation, while FINALLY achieves the highest perceptual quality yet lacks signal fidelity, a trend confirmed in Table 2. Diffusion-based methods yield more balanced results by directly operating in the waveform or complex STFT. TF-Locoformer preserves signal-level fidelity but suffers from residual perceptual artifacts and failure to recover lost details and naturalness (MOS, LSD). In contrast, TF-Restormer provides balanced improvements for signal fidelity and perceptual quality, with its streaming variant maintaining competitive effectiveness under causal constraints. This indicates its robustness across diverse degradations in a universal restoration setting. Note that all the compared models are offline methods.
@@ 5.2 Validation of stability across diverse scenarios
Model Signal fidelity Perceptual quality
PESQ SDR LSD MCD sBERT UTMOS DNSMOS
Input 1.98  8.56 1.27 5.40 0.91 2.90 2.45
Ground Truth 4.50 $\infty$ 0.00 0.00 1.00 4.07 3.16
DB-AIAT 3.27 21.30 0.90 1.77 0.95 3.83 3.13
MP-SENet 3.61 21.03 0.85 1.58 0.95 3.86 3.12
TF-Locoformer 3.30 23.82 0.92 3.58 0.95 3.93 3.20
VoiceFixer 2.40 -1.12 0.97 7.40 0.90 3.50 3.08
UNIVERSE 2.84 18.77 1.17 2.20 0.92 3.75 3.03
FINALLY 2.94   4.60 - - - 4.32 3.22
TF-Restormer (off) 3.41 19.45 0.75 1.54 0.95 4.14 3.14
@@
TF-Restormer (on) 2.89 16.43 0.85 2.16 0.93 4.05 3.09
Next, we evaluate TF-Restormer on the VCTK+DEMAND focusing on denoising. In Table 2, we compare against DB-AIAT (Yu et al., 2022), MP-SENet (Lu et al., 2023), and TF-Locoformer as dedicated denoising models, and VoiceFixer, UNIVERSE, and FINALLY as universal restoration baselines. Since the input speech is already well preserved and only corrupted by additive noise, it favors models that minimize unnecessary generation and faithfully retain the input signal. Accordingly, dedicated denoising models outperform universal restoration models in terms of signal fidelity, as they are optimized to suppress noise without altering intact regions. In contrast, restoration models risk degrading reliability by over-modifying clean inputs, making them less trustworthy for such simple cases. While not surpassing dedicated denoising models in raw signal metrics, TF-Restormer achieves more consistent gains, showing strong generalization despite being designed for universal restoration. We additionally include a dedicated TF-Restormer variant; as expected, this task-matched version achieves the higher signal-fidelity scores.
@@
Method 8 $\rightarrow$ 16kHz 8 $\rightarrow$ 24kHz 8 $\rightarrow$ 44.1kHz 16 $\rightarrow$ 48kHz
LSD NISQA LSD NISQA LSD NISQA LSD NISQA
clean (SSR only)
Input 2.53 3.78 2.91 3.78 3.44 3.78 3.17 4.40
@@
TF-Restormer (off) 0.89 4.53 0.95 4.61 1.01 4.54 0.97 4.62
@@
Noisy-distorted (GSR + SSR)
Input 3.36 1.91 3.49 1.91 3.64 1.91 3.48 1.73
@@
StoRM 1.76 3.97 - - - - - -
UNIVERSE++ 1.79 3.39 - - - - - -
TF-Restormer (off) 1.16 4.49 1.21 4.54 1.18 4.52 1.18 4.54
TF-Restormer (on) 1.30 4.42 1.31 4.49 1.30 4.46 1.26 4.46
Finally, we experiment on the SSR task in Table 3, using a single model that directly supports arbitrary output sampling rates. For clean cases, we compare against dedicated super-resolution models: NVSR (Liu et al., 2022a), Frepainter (Kim et al., 2024), and AP-BWE (Lu et al., 2025), as well as VoiceFixer as a universal restoration baseline. As in Table 2, since the low-band of the input speech remains intact, dedicated models that concentrate on reconstructing the upper bands are favored. Unlike conventional approaches that rely on fixed input–output rates and often require zero-padding or redundant resampling, TF-Restormer leverages extension queries to dynamically expand the spectrum. With this versatility, TF-Restormer shows stable performance comparable to the dedicated models, faithfully retaining clean low-frequency regions while effectively generating high-frequency components. When the training is optimally aligned with the conventional method, the dedicated version of the proposed model shows improved results. In addition, under noisy-distorted conditions, TF-Restormer simultaneously restores corrupted regions and reconstructs missing high bands, demonstrating robust generalization beyond pure super-resolution. Overall, these results suggest the advantage of our model as a universal restoration framework that achieves bandwidth extension without sacrificing signal fidelity or requiring explicit resampling.
Table 4: Evaluation of non-intrusive MOS results on real-recorded data from VoxCeleb and URGENT 2025 blind testset.
Model UTMOS DNSMOS
Input 2.76 2.72
VoiceFixer 2.60 3.08
DEMUCS 3.51 3.27
StoRM 3.29 3.17
HiFi-GAN-2 3.67 3.32
FINALLY 4.05 3.31
TF-Restormer 3.98 3.34
(a) VoxCeleb
Model UTMOS NISQA DNSMOS
Input 1.55 1.58 1.90
Bobbsun(R.1) 2.09 3.22 2.88
rc(R.2) 2.03 2.92 2.83
Xiaobin(R.3) 2.16 3.24 2.92
wataru9871(R.13) 2.53 3.74 3.10
LLaSE-G1 2.09 2.93 2.80
UniSE 2.85 3.72 3.17
TF-Restormer 3.37 4.37 3.13
(b) URGENT 2025 blind test set
5.3 MOS evaluation on additional datasets
VoxCeleb
We evaluate TF-Restormer on 50 real-recorded utterances (Su et al., 2020) from VoxCeleb1 (Nagrani et al., 2017) and compared with conventional method including DEMUCS (Défossez et al., 2019) and HiFi-GAN-2 (Su et al., 2021). As shown in Table 4(a), TF-Restormer achieves perceptual MOS scores (UTMOS, DNSMOS) comparable to recent vocoder- and diffusion-based models. While FINALLY (Babaev et al., 2024) remains one of the strongest perceptual-quality systems, our unified architecture delivers similarly natural outputs despite not being specialized for perceptual enhancement, demonstrating competitive robustness on real speech recordings.
URGENT 2025
Finally, we report non-intrusive MOS metrics on the URGENT 2025 blind test set (Saijo et al., 2025) compared to participating teams and lastest models including LLaSE-G1 (Kang et al., 2025) and UniSE (Yan et al., 2025). The official ranking incorporates both non-intrusive and intrusive measures, and the latter tend to favor deterministic bandwidth-preserving approaches while penalizing generative or reconstructive models. As a result, the top-ranked systems (Sun et al., 2025; Chao et al., 2025; Rong et al., 2025) are predominantly deterministic enhancers and generally obtain lower perceptual MOS. In contrast, TF-Restormer achieves natural-sounding outputs with strong non-intrusive MOS scores, showing that the model maintains stable perceptual quality on the URGENT blind test set as well.
Table 5: Ablation study on encoder-decoder design. VCTK-ND denote noisy-distorted input from VCTK data in Table 3.
Case Size(M) MAC(G) LSD NISQA
VCTK (SSR, 8 $\to$ 16kHz)
encoder-only 11.6 151.3 2.12 4.21
encoder-decoder w/o MHCA 30.8 252.4 1.04 4.48
encoder-decoder w/ MHCA 30.1 240.8 0.89 4.53
encoder-decoder w/ MHCA(small) 10.9 89.2 1.36 4.38
VCTK (SSR, 8 $\to$ 44.1kHz)
encoder-only 11.6 415.1 3.25 3.49
encoder-decoder w/o MHCA 30.8 340.4 1.35 4.26
encoder-decoder w/ MHCA 30.1 308.4 1.01 4.54
encoder-decoder w/ MHCA(small) 10.9 156.8 1.44 4.21
VCTK-ND (SSR+GSR, 8 $\to$ 16kHz)
encoder-only 11.6 151.3 2.23 3.72
encoder-decoder w/o MHCA 30.8 252.4 1.20 4.33
encoder-decoder w/ MHCA 30.1 240.8 1.16 4.54
encoder-decoder w/ MHCA(small) 10.9 89.2 1.48 4.20